OpenStudy (anonymous):

Find the probability of drawing a pink card, not replacing it, and then drawing another pink card. 2 blue 3 green 1 yellow 4 pink

OpenStudy (amistre64):

4/10 3/9 ---- 12/90 ----- = 4/30 = 2/15 maybe 3

OpenStudy (saifoo.khan):

hey, that blue card is lost.. Now?